OpenStudy (anonymous):

What is the solution of log3x + 4125 = 3? x = 1/3 x = 1 x = 7/3 x = 4

OpenStudy (chrisdbest):

x=4

OpenStudy (chrisdbest):

i think

OpenStudy (chrisdbest):

think about it this way, plug in all values for x and c wat anwer u get

OpenStudy (mathstudent55):

What exactly is the equation? Is the 3 the base of the log? Is it \(\log_3 x + 4125 = 3\) ? Is it \(\log (3 x) + 4125 = 3\) ? None of the choices work for either of the two equations I wrote. I don't understand what your equation is. Can you draw it or use the equation editor?

OpenStudy (anonymous):

I would put x = 1/3...but maybe I am wrong....